To remove xy term from the second degree equation 5x2+8xy+5y2+3x+2y+5=0,the coordinates axes are rotated through an angle∆,then ∆ equals to

Grade:11

1 Answers

Arun
3 years ago
Dear Priti
 
theta = 1/2 {tan-1 2h/(a-b)}       
         =1/2 {tan-1 8/5-5 }
         =1/2 {tan-18/0}
         =1/2 {pie/2}
         =pie/4
 
Hope it helps
 
Thanks and regards